Find the exponential decay function for a radioactive material that has a half-life of 1234 years. How long will it take until only 10% of the material remains?
Answer by stanbon(48569)

Find the exponential decay function for a radioactive material that has a half-life of 1234 years. How long will it take until only 10% of the material remains?
----------
A(t) = A(o)(1/2)^(t/1234)
-----------
0.10A(o) = A(o)(1/2)^(t/1234)
0.10 = (1/2)^(t/1234)
Take the log of both sides to get:
log0.10 = (t/1234)log(1/2)
t/1234 = [log0.10]/[log(1/2) = 3.321928095...
t = 4099.259269 years
===============
Cheers,
Stan H.